OpenStudy (anonymous):

Prove that |cos(x) - 1| ≤ |x| for all x. (Use Mean Value Theorem)

OpenStudy (anonymous):

I'll just outline the idea of the proof: first, there are definitely some values of x for which the inequality holds (e.g. pi/2) now suppose there is a point x* such that the inequality is not true: that implies that the LHS of the inequality goes from below the y=x line to above the y=x line. (using the language of mean value theorem, this means the secant line drawn from x=pi/2 to x = x* has a slope greater than 1). Now mean value theorem would say that there exists a point c such that f'(c) = slope of secant line > 1. look at your function f(x) = cos(x) - 1, its derivative is ~sin(x), which is bound by -1 and 1. and there's the contradiction that completes the proof.

OpenStudy (anonymous):

Thanks for the detailed response cnknd!